OpenStudy (anonymous):

Tough question: If the numerator of a fraction is increased by 6, the value of the fraction will increase by 1. If the denominator of the original fraction is increased by 36, the value of the original fraction will decrease by 1. What is the fraction?

OpenStudy (anonymous):

Wait a sec, I think i got it.

OpenStudy (anonymous):

I got it! 7/6

OpenStudy (anonymous):

you have to solve three equations for three variables. number 1: A=x/y; x is the numerator, y is the denominator, A is the value of the fraction so therefore from what we are given, number 2: A+1=(x+6)/y and number 3: A-1=x/(y+36) if you solve those three equations by substitution you will attain a value for each x and y, being the values of the original fraction x/y

OpenStudy (anonymous):

I made a systems that I didn't think would work: (x+6)/y = 1 + x/y x/(y + 36) = x/y - 1